OpenStudy (anonymous):

the probability that shania is on time for school is 1/2. find the probability that shania arrives on time for school for the next 5 days. express your answer as a percent,to the nearest tenth of a percent?

OpenStudy (anonymous):

\[(\frac{1}{2})^5=\frac{1}{2^5}=\frac{1}{32}\] not very likely is it?

OpenStudy (anonymous):

not sure i understand

OpenStudy (anonymous):

each day there is a 50% chance of getting to school on time i.e. probability is \(\frac{1}{2}\) for each day now you have 5 days in a row you want this to happen, so you multiply the probability by itself 5 times, better known as \(\left(\frac{1}{2}\right)^5\)

OpenStudy (anonymous):

so let me see if i got this right i am going to multiply 1times 5 and 2 times 5